OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

3x-y=-3 -y=-3-3x y=3+3x parallel means the slope is the same. Therefore, m=3 (4,-2) Plug the values we have into y=mx+b (-2)=(3)(4)+b -2=12+b b=-14 so... y=3x-14

OpenStudy (anonymous):

thank you ! will you help on another ?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

ok so you do the same thing.. y=-3x+12 m=-3 (-1,6) so... (6)=(-3)(-1)+b 6=3+b b=3 and... y=-3x+3

OpenStudy (anonymous):

thank you thank you so much! Last one ? lol

OpenStudy (anonymous):

haha sure!

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (anonymous):

ok so perpendicular slopes are the negative reciprical of the original slope. In this problem, the original slope is 3/2. Therefore the new slope or m is going to be -2/3. so... m=-2/3 (-6,3) (3)=(-2/3)(-6)+b 3=4+b b=-1 so... y=-2/3x-1

OpenStudy (anonymous):

Your the best! @kasper15

OpenStudy (anonymous):

Thanks!